# Radial Transport in High-Redshift Disk Galaxies Dominated by Inflowing Streams

### v1

@@ 5.1.1 Metallicity-Based Selection
Streams, due to their recent accretion onto the galaxy, are likely to have relatively low metallicities. Therefore, in a crude way, we can use a threshold metallicity, $Z_{\rm cut}$, to separate streams from the rest of the cold gas. To determine this metallicity threshold for a particular snapshot, we first find the mass-weighted cold gas metallicity distribution outside the disk radius, between $r=R_{\rm d}$ to $2R_{\rm d}$ and $z=-H_{\rm d}$ to $H_{\rm d}$, and then select the $99^{\rm th}$ percentile of this distribution as $Z_{\rm cut}$. Cold gas with $Z<Z_{\rm cut}$ at any $r$ is classified as streams and the rest as non-streams. The inherent assumption made in this selection procedure is that almost all the cold gas beyond $R_{\rm d}$ is in the form of recently accreted streams. Therefore, the maximum metallicity of this gas is the maximum metallicity of streams and can be used as a threshold to distinguish between streams and non-streams within the disk. Note that here we ignore the possibility that the gas beyond $R_{\rm d}$ may also have a metal-enriched, outflowing cold component ejected by stellar winds and supernovae (e.g. Veilleux et al., 2020), which can potentially bias this selection. However, the inflowing cold streams may themselves have higher metallicities because of the entrainment of initially hot, recycled, enriched gas that has been previously ejected from the galaxy (Strawn et al., 2021; Aung et al., 2024).
To give an example, in Fig. 6, the left-hand panel shows the 2D metallicity map, $Z_{\rm 2D}$, for VELA 7 at $z=1.5$ in a face-on projection of dimensions $4R_{\rm d}\times 4R_{\rm d}$, centered on the galactic center, obtained after mass-weighted averaging the metallicites ($Z$) of the cold gas cells along the $z$-axis between $-H_{\rm d}$ to $H_{\rm d}$. The red and cyan, dashed circles have radii equal to $R_{\rm d}$ and $2R_{\rm d}$, respectively. The image shows an overall negative metallicity gradient from small to large $r$, i.e., metallicity decreases as the cylindrical distance from the galactic center increases. This is consistent with the assumption that the cold gas outside the disk ($r\gtrsim R_{\rm d}$) likely consists almost entirely of low-metallicity streams, whereas that within the disk ($r\lesssim R_{\rm d}$) is a mix of both streams and non-streams. In the right-hand panel, the mass-weighted histogram of $Z$ for all cold gas cells between $r=R_{\rm d}$ to $2R_{\rm d}$ and $z=-H_{\rm d}$ to $H_{\rm d}$ is shown, normalized by the total cold gas mass inside this region. The black, dashed vertical line at $Z=0.009$ indicates $Z_{\rm cut}$, the $99^{\rm th}$ percentile of this distribution, and the threshold metallicity that separates streams from non-streams.
@@ 5.1.2 Streamline-Based Selection
Alternatively, we try to identify streams by tracing the path of each cold gas cell back in time. For simplicity, we assume that the gas velocity field remains stationary during the time over which we trace back the flow. Therefore, pathlines are the same as instantaneous streamlines, and cold gas cells in a given snapshot can be traced using the gas velocity data of that snapshot only. The tracing time, $t_{\rm trace}$, is adopted to be the average disk dynamical time, $t_{\rm trace}=R_{\rm d}/\langle V_{\rm rot}\rangle$. Starting from its current 3D position, ${r_{t}}$, at time, $t$, a cold gas cell in a given snapshot is traced back in time along its pathline (i.e., the streamline that originates from this position) to yield its initial 3D position, ${r_{i}}$, at time $t-t_{\rm trace}$. As streams are recently accreted material, the cells which move inwards over $t_{\rm trace}$, i.e., |${r_{i}}|>|\@vec{r_{t}}|$ |, and by sufficiently large distances, i.e., $d=|\@vec{r_{t}}-\@vec{r_{i}}|>d_{\rm cut}$, are classified as streams. The rest are classified as non-streams. By trial and error, we choose $d_{\rm cut}=0.1R_{\rm d}$, such that most cold gas cells that visually appear to be a part of streams get correctly identified by this procedure. Changing $d_{\rm cut}$ by a factor of two up or down has no significant impact on our results.
As an example, for VELA 7 at $z=1.5$, the pathlines of a random selection of cold gas cells within the disk, traced back in time over the interval $t_{\rm trace}$ and projected onto the $x$-$y$ plane, are shown in Fig. 7, separated into streams (left-hand panel) and non-streams (right-hand panel). The pathlines of all cold gas cells within the disk are not shown for better clarity. The lines are color-coded by ±$\pm d$, in units of $R_{\rm d}$, with positive values indicating the cases where |${r_{i}}|>|\@vec{r_{t}}|$ | and negative values showing the opposite. The red circles denote the radial extent of the disk. We note that most of the pathlines in the left-hand panel can be divided into two distinct groups, both having large positive $d$ values, indicating the inflow of cold gas either from the top or bottom-right of the image, which then spiral onto the galaxy counter-clockwise. The two groups also broadly coincide in space with the regions of the galaxy that can be visually identified as two distinct streams in the radial velocity and surface density maps (see the top left- and right-hand panels of Fig. 2). On the other hand, the pathlines in the right-hand panel either have $d\approx 0$ (greenish lines), indicating little or no displacement over $t_{\rm trace}$, or have large negative $d$ values, indicating outflowing cold gas.
@@ 5.2 Stream and Non-Stream Signal: An Example
Figure 10: Stream and Non-Stream Radial Profiles: The left-, middle, and right-hand panels, respectively, show the radial mass flux, $F_{r}$ (in units of $\mathscr{M}/t_{\rm dyn}$), average radial velocity, $V_{r}$ (in units of $V_{\rm rot}$), and mass, $M$ (in units of $\mathscr{M}$), as functions of $r$ (in units of $R_{\rm d}$) for all cold gas (black), streams (blue), and non-streams (red) for VELA 7 at $z=1.5$. Here, $V_{\rm rot}$ and $\mathscr{M}$ are the average rotational velocity and mass of all cold gas at radius $r$, and $t_{\rm dyn}=r/V_{\rm rot}$ is the local dynamical time (see §2.2). The results from the metallicity-based selection are shown in the top row, while those from the streamline-based selection are plotted in the bottom row. The total cold gas radial mass flux is dominated by the streams at large $r$ (outside $\sim\!0.7\ R_{\rm d}$) and non-streams at small $r$ (inside $\sim\!0.3\ R_{\rm d}$) in both selections. In the metallicity-based selection, the flux dominance of streams at large $r$ stems from their corresponding mass dominance. In the streamline-based selection, however, the stream and non-stream masses are comparable at large $r$. Nevertheless, the non-stream flux is still small, as outflowing and inflowing non-stream fluxes of comparable magnitude almost cancel each other (see Fig 9), and the stream flux dominates.
@@
Figure 11: Stream and Non-Stream Statistics: The top, middle, and bottom rows, respectively, show the radial mass flux, $F_{r}$ (in units of $\mathscr{M}/t_{\rm dyn}$), average radial velocity, $V_{r}$ (in units of $V_{\rm rot}$), and mass, $M$ (in units of $\mathscr{M}$) as functions of $r$ (in units of $R_{\rm d}$) for streams (left-hand panels), non-streams (middle panels), and all cold gas (right-hand panels), with the solid lines indicating the medians over the disk galaxy sample and the envelopes representing the $16^{\rm th}-84^{\rm th}$ percentiles. The results for the metallicity-based selection are shown in blue, while those for the streamline-based selection are plotted in green. In both selections, the median cold gas flux at large $r$ ($r\gtrsim 0.5R_{\rm d}$ in the metallicity-based selection and $r\gtrsim 0.3R_{\rm d}$ in the streamline-based selection) is dominated by the streams. At smaller $r$, the median stream flux tends to zero and is less than the median non-stream flux, so the median cold gas flux, although very small, is dominated by the non-streams. The median stream $V_{r}$ is almost always negative, with typically larger magnitudes than the median non-stream $V_{r}$, especially at large $r$. Thus, we conclude that, on average, the radial transport of cold gas in the VELA disk galaxy sample at large $r$ is dominated by the inflowing motion of recently accreted streams.
@@
Figure 12: Positive vs Negative Flux: The top and bottom rows, respectively, show the radial mass flux, $F_{r}$ (in units of $\mathscr{M}/t_{\rm dyn}$), from the inflowing and outflowing regions of streams (left-hand panels), non-streams (middle panels), and all cold gas (right-hand panels) as a function of $r$ (in units of $R_{\rm d}$), with the solid lines indicating the medians over the disk galaxy sample and the envelopes representing the $16^{\rm th}-84^{\rm th}$ percentiles. The results from the metallicity- and streamline-based selections are shown in blue and green, respectively. Note that in the streamline-based selection, inflowing and outflowing non-stream fluxes of comparable magnitude (which are also similar in magnitude to the inflowing stream flux) almost cancel each other, yielding a total non-stream flux that is close to zero at the median level, albeit with a large scatter. In the metallicity-based selection, inflowing and outflowing non-stream fluxes are also of comparable magnitude and can cancel each other. However, even by themselves, they are negligible compared to the stream flux, except at small $r$ (inside $\sim\!0.5\ R_{\rm d}$), where the stream flux is also small. As such, the total non-stream flux in this selection is also close to zero, with very little scatter.
@@
In Figs. 8 and 9, we compare the face-on radial velocity (top rows) and surface density (bottom rows) maps of streams and non-streams, respectively, obtained with the metallicity- (middle panels) and streamline-based (right-hand panels) selections for VELA 7 at $z=1.5$. These maps are generated following the procedure outlined in §3.1, except that instead of all cold gas cells, only cells identified as streams or non-streams are used, as relevant. To check the results of the two selections against what visually appears as streams, the face-on radial velocity and surface density maps of all cold gas in a slightly larger region of dimensions $3R_{\rm d}\times 3R_{\rm d}$, as shown in the top right- and left-hand panels of Fig. 2, are re-plotted here in the top and bottom left-hand panels, respectively. Note that the actual classification of cold gas cells into streams and non-streams is only done within the disk (i.e., within a cylinder of radius $R_{\rm d}$ and height $H_{\rm d}$), as shown in the middle and right-hand panels. The red, dashed circles indicate the radial extent of the disk.
@@
For the metallicity-based selection (top row), the mass in streams (non-streams) increases (decreases) with an increase in $r$, attaining values close to $\mathscr{M}$ (zero) near $r=R_{\rm d}$. Consequently, the cold gas radial mass flux, which is a sum of the stream and non-stream flux, is also dominated by the streams at large $r$, outside $\sim\!0.7\ R_{\rm d}$. Between $\sim\!0.3\ R_{\rm d}$ to $\sim\!0.7\ R_{\rm d}$, the stream and non-stream fluxes are comparable in magnitude, with the non-stream flux oscillating between positive and negative values, while the stream flux is always negative. Inside $\sim\!0.3\ R_{\rm d}$, the stream flux tends to zero, and the cold gas flux, albeit very small, is dominated by the non-streams and directed outwards. The stream $V_{r}$ is mostly negative (except within $\sim\!0.3\ R_{\rm d}$), reaching values close to $-0.2\ V_{\rm rot}$ near $r=R_{\rm d}$. The non-stream $V_{r}$ oscillates between positive and negative values. The cold gas $V_{r}$, which is a mass-weighted average of the stream and non-stream $V_{r}$, is almost the same as the stream $V_{r}$ outside $\sim\!0.7\ R_{\rm d}$ and as the non-stream $V_{r}$ inside $\sim\!0.3\ R_{\rm d}$, owing to the mass dominance of the streams and non-streams in these two regions, respectively.
@@
For the streamline-based selection (bottom row), the mass in streams (non-streams) increases (decreases) with $r$ only up to about $\sim\!0.4\ R_{\rm d}$, after which it roughly stays the same up to $\sim\!0.7\ R_{\rm d}$, and then decreases (increases). The cold gas radial mass flux outside $\sim\!0.7\ R_{\rm d}$ is dominated by the streams, as the non-stream flux is much smaller in magnitude. However, note that this is because of outflowing and inflowing non-stream fluxes of comparable magnitude almost canceling each other (see right-hand panels of Fig. 9; contributions to the total non-stream flux from regions b and c almost cancel each other) and not because of small non-stream mass as in the metallicity-based selection. Similar to the metallicity-based selection, between $\sim\!0.4\ R_{\rm d}$ to $\sim\!0.7\ R_{\rm d}$, the stream and non-stream fluxes are comparable in magnitude, with the non-stream flux being mostly positive and the stream flux being negative. Between $\sim\!0.3\ R_{\rm d}$ to $\sim\!0.4\ R_{\rm d}$, however, the stream flux again dominates, but inside $\sim\!0.3\ R_{\rm d}$, the stream flux tends to zero, and the cold gas flux, albeit very small, is dominated by the non-streams as in the metallicity-based selection. The stream $V_{r}$ is mostly negative, reaching values close to $-0.4\ V_{\rm rot}$ near $r=R_{\rm d}$, while the non-stream $V_{r}$ is mostly positive and smaller in magnitude than the stream $V_{r}$. As the stream and non-stream masses are comparable outside $\sim\!0.4\ R_{\rm d}$, the cold gas $V_{r}$, being a mass-weighted average of the stream and non-stream $V_{r}$, is somewhere in between (and negative). Inside $\sim\!0.4\ R_{\rm d}$, the non-stream mass dominates, and the cold gas $V_{r}$ is close to the non-stream $V_{r}$.

### v2

@@ 5.2.1 Metallicity-Based Selection
Streams, due to their recent accretion onto the galaxy, are likely to have relatively low metallicities. Therefore, in a crude way, we can use a threshold metallicity, $Z_{\rm cut}$, to separate streams from the rest of the cold gas. To determine this metallicity threshold for a particular snapshot, we first find the mass-weighted cold gas metallicity distribution outside the disk radius, between $r=R_{\rm d}$ to $2R_{\rm d}$ and $z=-H_{\rm d}$ to $H_{\rm d}$, and then select the $99^{\rm th}$ percentile of this distribution as $Z_{\rm cut}$. Our results are robust against the particular choice for this threshold, as adopting $Z_{\rm cut}$ to be the $90^{\rm th}$ or $95^{\rm th}$ percentile instead of the $99^{\rm th}$ has no significant impact (see Appendix B). Cold gas with $Z<Z_{\rm cut}$ at any $r$ is classified as streams and the rest as non-streams. The inherent assumption made in this selection procedure is that almost all the cold gas beyond $R_{\rm d}$ is in the form of recently accreted streams. Therefore, the maximum metallicity of this gas is the maximum metallicity of streams and can be used as a threshold to distinguish between streams and non-streams within the disk. Note that here we ignore the possibility that the gas beyond $R_{\rm d}$ may also have a metal-enriched, outflowing cold component ejected by stellar winds and supernovae (e.g. Veilleux et al., 2020), which can potentially bias this selection. However, the inflowing cold streams may themselves have higher metallicities because of the entrainment of initially hot, recycled, enriched gas that has been previously ejected from the galaxy (Strawn et al., 2021; Aung et al., 2024).
To give an example, in Fig. 7, the left-hand panel shows the 2D metallicity map, $Z_{\rm 2D}$, for VELA 7 at $z=1.5$ in a face-on projection of dimensions $4R_{\rm d}\times 4R_{\rm d}$, centered on the galactic center, obtained after mass-weighted averaging the metallicites ($Z$) of the cold gas cells along the $z$-axis between $H_{\rm d}$ to $H_{\rm d}$. The red and cyan, dashed circles have radii equal to $R_{\rm d}$ and $2R_{\rm d}$, respectively. The image shows an overall negative metallicity gradient from small to large $r$, i.e., metallicity decreases as the cylindrical distance from the galactic center increases. This is consistent with the assumption that the cold gas outside the disk ($r\gtrsim R_{\rm d}$) likely consists almost entirely of low-metallicity streams, whereas that within the disk ($r\lesssim R_{\rm d}$) is a mix of both streams and non-streams. In the right-hand panel, the mass-weighted histogram of $Z$ for all cold gas cells between $r=R_{\rm d}$ to $2R_{\rm d}$ and $z=-H_{\rm d}$ to $H_{\rm d}$ is shown, normalized by the total cold gas mass inside this region. The black, dashed vertical line at $Z=0.009$ indicates $Z_{\rm cut}$, the $99^{\rm th}$ percentile of this distribution, and the threshold metallicity that separates streams from non-streams.
@@ 5.2.2 Streamline-Based Selection
Alternatively, we try to identify streams by tracing the path of each cold gas cell back in time. For simplicity, we assume that the gas velocity field remains stationary during the time over which we trace back the flow. Therefore, pathlines are the same as instantaneous streamlines, and cold gas cells in a given snapshot can be traced using the gas velocity data of that snapshot only. The tracing time, $t_{\rm trace}$, is adopted to be the average disk dynamical time, $t_{\rm trace}=R_{\rm d}/\langle V_{\rm rot}\rangle$. Starting from its current 3D position, $\@vec{r_{t}}$, at time, $t$, a cold gas cell in a given snapshot is traced back in time along its pathline (i.e., the streamline that originates from this position) to yield its initial 3D position, $\@vec{r_{i}}$, at time $t-t_{\rm trace}$. As streams are recently accreted material, the cells which move inwards over $t_{\rm trace}$, i.e., $|\@vec{r_{i}}|>|\@vec{r_{t}}|$, and by sufficiently large distances, i.e., $d=|\@vec{r_{t}}-\@vec{r_{i}}|>d_{\rm cut}$, are classified as streams. The rest are classified as non-streams. By trial and error, we choose $d_{\rm cut}=0.1R_{\rm d}$, such that most cold gas cells that visually appear to be a part of streams get correctly identified by this procedure. Changing $d_{\rm cut}$ by a factor of two up or down has no significant impact on our results (see Appendix B). We note that the streamline-based selection is biased towards detecting only inflowing material as streams, while the metallicity-based selection doesn’t have this bias.
As an example, for VELA 7 at $z=1.5$, the pathlines of a random selection of cold gas cells within the disk, traced back in time over the interval $t_{\rm trace}$ and projected onto the $x$-$y$ plane, are shown in Fig. 8, separated into streams (left-hand panel) and non-streams (right-hand panel). The pathlines of all cold gas cells within the disk are not shown for better clarity. The lines are color-coded by $\pm d$, in units of $R_{\rm d}$, with positive values indicating the cases where $|\@vec{r_{i}}|>|\@vec{r_{t}}|$ and negative values showing the opposite. The red circles denote the radial extent of the disk. We note that most of the pathlines in the left-hand panel can be divided into two distinct groups, both having large positive $d$ values, indicating the inflow of cold gas either from the top or bottom-right of the image, which then spiral onto the galaxy counter-clockwise. The two groups also broadly coincide in space with the regions of the galaxy that can be visually identified as two distinct streams in the radial velocity and surface density maps (see the top left- and right-hand panels of Fig. 2). On the other hand, the pathlines in the right-hand panel either have $d\approx 0$ (greenish lines), indicating little or no displacement over $t_{\rm trace}$, or have large negative $d$ values, indicating outflowing cold gas.
@@ 5.3 Stream and Non-Stream Signal: An Example
Figure 11: Stream and Non-Stream Radial Profiles: The left-, middle, and right-hand panels respectively show the radial mass flux ($F_{r}$, in units of $\mathscr{M}/t_{\rm dyn}$), average radial velocity ($V_{r}$, in units of $V_{\rm rot}$), and mass ($M$, in units of $\mathscr{M}$), as functions of radius $r$, normalized by $R_{\rm d}$, for all cold gas (black), streams (blue), and non-streams (red) in VELA 7 at $z=1.5$. The top row presents results from the metallicity-based classification, and the bottom row corresponds to the streamline-based classification. In both cases, the total radial mass flux of cold gas is dominated by streams at large radii (beyond $\sim\!0.7\ R_{\rm d}$) and by non-streams at small radii (within $\sim\!0.3\ R_{\rm d}$). In the metallicity-based selection, the dominance of stream flux at large $r$ primarily reflects the higher mass fraction of streams in that region. In contrast, the streamline-based method shows roughly equal masses for streams and non-streams at large $r$; however, the non-stream mass flux remains low. This is because the inflowing and outflowing contributions from non-stream gas nearly cancel out (see Fig. 10), leaving streams as the main net contributor to the total radial mass flux.
@@
Figure 12: Stream and Non-Stream Statistics: The top, middle, and bottom rows display, respectively, the radial mass flux ($F_{r}$, in units of $\mathscr{M}/t_{\rm dyn}$), average radial velocity ($V_{r}$, in units of $V_{\rm rot}$), and mass ($M$, in units of $\mathscr{M}$), as functions of radius $r$, normalized by $R_{\rm d}$, for streams (left panels), non-streams (middle panels), and all cold gas (right panels). Solid lines represent the median values across the disk galaxy sample, while shaded regions indicate the $16^{\rm th}$$84^{\rm th}$ percentile ranges. The results from the metallicity-based selection are shown in blue, and those from the streamline-based selection are shown in green. In both classifications, the median cold gas mass flux at large radii is dominated by streams — beyond $r\gtrsim 0.5R_{\rm d}$ in the metallicity-based selection and beyond $r\gtrsim 0.3R_{\rm d}$ in the streamline-based selection. At smaller radii, the median stream flux approaches zero and falls below that of the non-stream component, resulting in a total cold gas flux that, while small, is primarily contributed by non-streams. The median radial velocity of streams remains predominantly negative across all radii, and its magnitude tends to exceed that of the non-streams, particularly at large $r$. Thus, on average, the radial transport of cold gas in the outer regions of the VELA disk galaxy sample is driven primarily by the infall of recently accreted streams.
@@
Figure 13: Positive vs Negative Flux: The top and bottom rows illustrate the radial mass flux ($F_{r}$, in units of $\mathscr{M}/t_{\rm dyn}$), from inflowing and outflowing regions, respectively, for streams (left-hand panels), non-streams (middle panels), and the total cold gas (right-hand panels), plotted as a function of radius $r$, normalized by $R_{\rm d}$. Solid lines represent the median values across the disk galaxy sample, and shaded regions correspond to the $16^{\rm th}$$84^{\rm th}$ percentile ranges. Results from the metallicity- and streamline-based classifications are shown in blue and green, respectively. In the streamline-based selection, inflowing and outflowing fluxes from non-stream gas are nearly equal in magnitude, comparable also to the inflowing stream flux, and largely cancel each other out. This leads to a total non-stream flux that is close to zero at the median level, though it exhibits significant scatter. A similar cancellation is also observed in the metallicity-based selection, where inflowing and outflowing non-stream fluxes are also comparable but individually much smaller than the stream flux, except at small radii ($r\lesssim 0.5\ R_{\rm d}$), where the stream flux diminishes. As a result, the total non-stream flux in the metallicity-based selection also approaches zero, but with minimal scatter.
@@
In Figs. 9 and 10, we compare the face-on radial velocity (top rows) and surface density (bottom rows) maps of streams and non-streams, respectively, obtained with the metallicity- (middle panels) and streamline-based (right-hand panels) selections for VELA 7 at $z=1.5$. These maps are generated following the procedure outlined in §3.1, except that instead of all cold gas cells, only cells identified as streams or non-streams are used, as relevant. To check the results of the two selections against what visually appears as streams, the face-on radial velocity and surface density maps of all cold gas in a slightly larger region of dimensions $3R_{\rm d}\times 3R_{\rm d}$, as shown in the top right- and left-hand panels of Fig. 2, are re-plotted here in the top and bottom left-hand panels, respectively. Note that the actual classification of cold gas cells into streams and non-streams is only done within the disk (i.e., within a cylinder of radius $R_{\rm d}$ and height $H_{\rm d}$), as shown in the middle and right-hand panels. The red, dashed circles indicate the radial extent of the disk.
@@
For the metallicity-based selection (top row), the mass in streams (non-streams) increases (decreases) with an increase in $r$, attaining values close to $\mathscr{M}$ (zero) near $r=R_{\rm d}$. Consequently, the cold gas radial mass flux, which is a sum of the stream and non-stream flux, is also dominated by the streams at large $r$, outside $\sim\!0.7\ R_{\rm d}$. Between $\sim\!0.3\ R_{\rm d}$ to $\sim\!0.7\ R_{\rm d}$, the stream and non-stream fluxes are comparable in magnitude, with the non-stream flux oscillating between positive and negative values, while the stream flux is always negative. Inside $\sim\!0.3\ R_{\rm d}$, the stream flux tends to zero, and the cold gas flux, albeit very small, is dominated by the non-streams and directed outwards. The stream $V_{r}$ is mostly negative (except within $\sim\!0.3\ R_{\rm d}$), reaching values close to $0.2\ V_{\rm rot}$ near $r=R_{\rm d}$. The non-stream $V_{r}$ oscillates between positive and negative values. The cold gas $V_{r}$, which is a mass-weighted average of the stream and non-stream $V_{r}$, is almost the same as the stream $V_{r}$ outside $\sim\!0.7\ R_{\rm d}$ and as the non-stream $V_{r}$ inside $\sim\!0.3\ R_{\rm d}$, owing to the mass dominance of the streams and non-streams in these two regions, respectively.
@@
For the streamline-based selection (bottom row), the mass in streams (non-streams) increases (decreases) with $r$ only up to about $\sim\!0.4\ R_{\rm d}$, after which it roughly stays the same up to $\sim\!0.7\ R_{\rm d}$, and then decreases (increases). The cold gas radial mass flux outside $\sim\!0.7\ R_{\rm d}$ is dominated by the streams, as the non-stream flux is much smaller in magnitude. However, note that this is because of outflowing and inflowing non-stream fluxes of comparable magnitude almost canceling each other (see right-hand panels of Fig. 10; contributions to the total non-stream flux from regions b and c almost cancel each other) and not because of small non-stream mass as in the metallicity-based selection. Similar to the metallicity-based selection, between $\sim\!0.4\ R_{\rm d}$ to $\sim\!0.7\ R_{\rm d}$, the stream and non-stream fluxes are comparable in magnitude, with the non-stream flux being mostly positive and the stream flux being negative. Between $\sim\!0.3\ R_{\rm d}$ to $\sim\!0.4\ R_{\rm d}$, however, the stream flux again dominates, but inside $\sim\!0.3\ R_{\rm d}$, the stream flux tends to zero, and the cold gas flux, albeit very small, is dominated by the non-streams as in the metallicity-based selection. The stream $V_{r}$ is mostly negative, reaching values close to $0.4\ V_{\rm rot}$ near $r=R_{\rm d}$, while the non-stream $V_{r}$ is mostly positive and smaller in magnitude than the stream $V_{r}$. As the stream and non-stream masses are comparable outside $\sim\!0.4\ R_{\rm d}$, the cold gas $V_{r}$, being a mass-weighted average of the stream and non-stream $V_{r}$, is somewhere in between (and negative). Inside $\sim\!0.4\ R_{\rm d}$, the non-stream mass dominates, and the cold gas $V_{r}$ is close to the non-stream $V_{r}$.